# SageAttention2 Technical Report: Accurate 4 Bit Attention for Plug-and-play Inference Acceleration

### v1

@@ 3.3 Smooth Q
  Preprocessing: $K=K-\mathrm{mean}(K)$, $\overrightarrow{Q}_{m}=\mathrm{mean}(Q)$, $Q=Q-\overrightarrow{Q}_{m}$, $\Delta S=\text{GEMV}(\overrightarrow{Q}_{m},K^{\top})$, $\overrightarrow{V}_{m}=\mathrm{mean}(V),V=V-\overrightarrow{V}_{m}$.
  Quantization: ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{Q},\hat{Q})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{Q}}}(Q)$ // INT4 Per-warp. ${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{K}}}(K)$. ${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{V}}}(V)$. // FP8 Per-channel.
@@
     Load $\hat{Q}_{i}$ and $\delta_{Q}[i*c_{w}:(i+1)*c_{w}]$ into an SM ;
@@
        Load $\hat{K}_{j}$, $\hat{V}_{j}$, and $\delta_{K}[j]$ into the SM ;
@@
        $S_{i}^{j}[st:st+c_{w}]=\mathrm{Matmul}(\hat{Q}_{i}[st:st+c_{w}],\hat{K}_{j}^{\top})\times\delta_{Q}[st+w]\times\delta_{K}[j]$ + $\Delta S_{j}$; // Paralleled by $c_{w}$ warps.
@@
        $O_{i}^{j}=\mathrm{diag}(e^{m_{i}^{j-1}-m_{i}^{j}})^{-1}O_{i}^{j-1}+$ $\mathrm{Matmul}((\widetilde{P}_{i}^{j}*448)\text{.to(FP8.e4m3)},V_{j})$ ;
@@
Smooth V Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
98.25% 0.1980 0.2387
99.75% 0.0406 0.0773
@@
Table 7: Error of the FP8 Matmul instruction of mma(f8f8f32).
Precision of Accumulated Value E8M13 E8M23
3.4 Smooth V
To avoid the drawback of SageAttention where FP16 Matmul with an FP16 accumulator for $PV$ is only effective on GPUs such as the RTX4090, we adopt the approach of quantizing $P$ and $V$ to FP8 to leverage the universal acceleration of FP8 Tensor cores. However, we discover that the accumulator for the mma(f32f8f8f32) instruction on the Ada architecture is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. Specifically, for mma(f32f8f8f32) instruction $C=AB+D$, where $A,B$ are tensors in FP8 data type and $C,D$ are tensors with FP32 data type, we initialize the $A,B$ to zero and vary $D$ to test the data type of the accumulator. As shown in Table 7, when $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ matches the result of the mma(f16f16f32) instruction. However, when $D$ is initialized with more than 13 mantissa bits, the error of $C$ corresponds to the difference between the results of mma(f32f16f16f32) and mma(f32f8f8f32). Consequently, matrix multiplication of matrices $\widetilde{P}$ and $V$, quantized to FP8, incurs a certain degree of precision loss compared to using an FP32 accumulator. To mitigate this precision loss as much as possible, we propose to smooth $V$:
@@ 3.4 Smooth V
As shown in Figure 6, this strategy enhances the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ consistently features channel-wise outliers that are exclusively positive or negative, typically ranging between 8 and 9. Consequently, the values of $\widetilde{P}V$ could be quite large. However, the floating-point number representation range is not uniform—it is denser near zero. Therefore, by subtracting the mean along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in higher representational precision. We believe such an explanation and strategy could address lots of issues [^1] https://github.com/triton-lang/triton/issues/4476, https://github.com/triton-lang/triton/issues/5065 reported by the community. Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$. This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$. In other words, this method decomposes $V$ into two parts: $\overrightarrow{V_{m}}$ and $V$. For $V$, it centers the values of each column around zero, which leads to the dot product result of a row from the quantized $\widetilde{P}$ matrix with a column from the quantized $V$ matrix being closer to zero. This makes the representation of FP22 more accurate. Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.
Table 6 shows the attention accuracy on real tensors sampled from CogvideoX with and without smoothing $V$. It demonstrates that smoothing V could improve the accuracy of SageAttention2 when quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
3.5 Quantization for Q, K, P, V
Quantization for $Q,K$. We propose $\psi_{Q}(Q)$ and $\psi_{K}(K)$ in granularity of per-warp. This is first because per-channel quantization is not feasible since the scale factors of the inner axis of $QK^{\top}$ cannot used to do dequantization (Xiao et al., 2023a). Second, as shown in Table 3 and Table 3, we compare the average and worst-case accuracies of INT4 quantization at per-token, per-warp, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX. Results indicate that the accuracy of per-warp quantization is very close to per-token and outperforms a lot than per-block and per-tensor. Furthermore, per-warp quantization incurs less dequantization overhead than per-token as discussed in Section 3.2.
Quantization for $\widetilde{P},V$. We choose FP8, specifically the E4M3 data type, for $\psi_{P}(\widetilde{P})$ and $\psi_{V}(V)$ for two reasons. First, most GPUs have tensor cores that support FP8 Matmul operations, which are twice as fast as those using FP16. Second, Table 5 and Table 5 show the average and worst accuracy of different data types used for $\widetilde{P},V$ using real $Q,K,V$ across all layers of CogvideoX. We can see that the accuracy of using E4M3 is very close to that of using FP16 and superior to that of E5M2 and INT8. We propose to use $\psi_{P}(\widetilde{P})$ in per-block and $\psi_{V}(V)$ in per-channel for three reasons. First, per-channel quantization for $\widetilde{P}$ and per-token quantization for $V$ are not viable because dequantization requires scale factors of outer axis. Second, $\widetilde{P}=\mathrm{exp}(S_{i}-\mathrm{rowmax}(S_{i}))$, where $S_{i}$ is the Matmul result of a block of $Q$ and $K^{\top}$, the max value in each row of $\widetilde{P}$ is 1. Hence, we can assign a single static scale $s=\frac{1}{448}$ to a block $\widetilde{P}$, whose accuracy equals per-token quantization. Third, per-channel quantization can address the channel-wised outlier of $V$.
Table 8: Two kernel implementations of SageAttention2.
     Kernel      $\psi_{Q}(Q)$, $\psi_{K}(K)$      $\psi_{P}(\widetilde{P})$, $\psi_{V}(V)$
     SageAttn2-4b      INT4, per-warp      FP8, per-block;  FP8, per-channel
     SageAttn2-8b      INT8, per-warp      FP8, per-block;  FP8, per-channel
Figure 7: Mean and standard deviation of $cossim*(1-L1)$) of SageAttn-4b in different layers and timesteps for different inputs in Llama3.1 and CogvideoX.
3.6 Adaptive Quantization over Layer and Timestep
Based on the discussion in Section 3.5, we implement two attention kernels (See Table 8) based on a choice: Using INT4 or INT8 quantization for $\psi_{Q}(Q)$ and $\psi_{K}(K)$. The speed of these kernels is in the order (SageAttn2-4b$>$SageAttn2-8b), but the accuracy order is opposite. Although SageAttn2-4b could perform adequately in lots of tasks, for instance, it shows very slight end-to-end metric loss of Lambda benchmark on Llama3.1, SageAttn2-4b falls short in some hard scenarios. Consequently, we first conduct a detailed analysis of the accuracy performance on various models using SageAttn2-4b and SageAttn2-8b. Then, we propose an effective and easy-to-use adaptive method.
Accuracy across layers and timesteps. Figure 7 displays the mean and standard deviation of the accuracy, calculated using $CosSim*(1-L1)$, for SageAttn-4b across different layers and timesteps of different inputs in Llama3.1 and CogvideoX. It can be observed that specific layers and timesteps exhibit specific ranges of errors.
Our method. Our adaptive strategy is detailed as follows: First, we assess the average accuracy of SageAttn2-4b in comparison to full precision attention for each (layer, timestep) combination within a model, using several random inputs. For models without a timestep dimension, such as language models, the evaluation is confined to the layer dimension. Second, we will sort the (layer, timestep) combinations in descending order based on their $CosSim*(1-L1)$ values. We then identify the smallest $\xi$of these combinations. Finally, for the identified (layer, timestep) combinations, we consistently apply SageAttn2-8b across all other prompts. We call such a strategy SageAttn-mix and will report the values of $\xi$we used for each model in Section 4.1.
@@ 3.6 Adaptive Quantization over Layer and Timestep
Figure 10: Speed comparison between SageAttention2 and baselines (RTX4090, headdim=256).
@@
Figure 13: Speed comparison between SageAttention2 and baselines (L20, headdim=256).
@@ 4.1 Setup
Models. We validate the effectiveness of SageAttention2 across a diverse set of representative models from language, image, and video generation. Specifically, we conduct experiments on five models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B) (Yang et al., 2024) and Open-Sora (Zheng et al., 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
Datasets. Text-to-text models are evaluated on four zero-shot tasks: WikiText (Merity et al., 2022) to assess the model’s prediction confidence, LAMBADA (Paperno et al., 2016) evaluate contextual understanding, MMLU (Hendrycks et al., 2020) for measuring knowledge across various subjects, and Longbench (Bai et al., 2023) for evaluating bilingual, multitask, and comprehensive assessment of long context understanding capabilities. Text-to-video models are evaluated using the open-sora (Zheng et al., 2024) prompt sets. Flux is assessed on MJHQ-30K (Li et al., 2024). TIMM is evaluated on on three image datasets: ImageNet (Deng et al., 2009), ImageNet-Sketch (Sketch) (Wang et al., 2019), and ImageNet-Rendition (ImageNet-r) (Hendrycks et al., 2021).
Metrics. For text-to-text models, we use perplexity (ppl.) (Jelinek et al., 1977) for WikiText, Accuracy (Acc.) for LAMBADA and MMLU, and Longbench score (Bai et al., 2023). For text-to-video models, following Zhao et al. (2024), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024) to measure the text-video alignment; (VQA-a) and (VQA-t) to assess the video aesthetic and technical quality, respectively; and Flow-score (FScore) for temporal consistency (Wu et al., 2023). For text-to-image models, generated images are compared with the images in MJHQ-30K dataset in three aspects: FID (Heusel et al., 2017) and sFID (Salimans et al., 2016) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2024) for human preference. For TIMM, we use Accuracy.
@@
Baselines. (1) SmoothAttn. Following Qserve (Lin et al., 2024), we apply smooth quant for matrixes $Q,K$ with smoothing factor $\alpha=0.5$. (2) HadmdAttn. Following Quarot (Ashkboos et al., 2024), we apply random Hadamard transformation for matrixes $Q,K$ before INT4 quantization.
Hyper-parameters of SageAttn-mix. We use $\xi$ of 50% for Llama2, 60% for CogvideoX, 25% for Open-Sora, 30% for Llama3.1 and GLM4, 0% for Flux, and TIMM.
Figure 14: Comparison examples from CogvideoX, prompts are sampled from open-sora prompt sets.
Table 9: Average accuracy across all layers using different smooth methods.
@@
Table 10: Worst accuracy across all layers using different smooth methods.
@@
Table 11: End-to-end metrics loss across text, image, and video generation models.
@@
HadmdAttn 6.706 0.865 0.355 -
SmoothAttn 6.690 0.871 0.395 -
SageAttn2-4b 6.018 0.886 0.436 -
SageAttn2-mix 5.883 0.883 0.431 -
@@
HadmdAttn 7.661 0.756 0.502 44.62
SmoothAttn 7.087 0.788 0.551 43.76
SageAttn2-4b 6.219 0.808 0.617 48.61
SageAttn2-mix 6.131 0.816 0.629 49.01
@@
HadmdAttn 7.932 0.435 0.676 46.27
SmoothAttn 8.841 0.442 0.599 43.10
SageAttn2-4b 7.341 0.435 0.732 49.06
SageAttn2-mix 7.303 0.434 0.737 49.77
@@
CogvideoX Full-Precision 0.1836 0.9975 77.605 75.360 3.006
HadmdAttn 0.1742 0.9877 36.028 23.786 0.550
SmoothAttn 0.1763 0.9870 37.444 42.184 0.792
SageAttn2-4b 0.1813 0.9969 77.276 75.147 2.070
SageAttn2-mix 0.1816 0.9976 75.686 78.600 2.884
Open-Sora Full-Precision 0.1831 0.9996 46.713 59.553 0.368
SageAttn2-4b 0.1821 0.9994 42.270 55.965 0.364
SageAttn2-mix 0.1814 0.9994 44.509 59.097 0.383
@@
        Flux         Full-Precision        11.303        17.603        32.603        0.9745
        HadmdAttn        11.163        17.693        32.592        0.9638
        SmoothAttn        10.941        18.098        32.582        0.9613
        SageAttn2-4b        10.563        17.052        32.631        0.9747
@@
SageAttn2-4b 84.67% 45.07% 59.11%

### v3

@@ 3.3 INT4 Per-thread Quantization
  Preprocessing: $K=K-\mathrm{mean}(K)$, $\overrightarrow{Q}_{m}=\mathrm{mean}(Q)$, $Q=Q-\overrightarrow{Q}_{m}$, $\Delta S=\text{GEMV}(\overrightarrow{Q}_{m},K^{\top})$.
  Quantization: ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{Q},\hat{Q})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{Q}}}(Q)$, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{K}}}(K)$. //per-thread. ${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{V}}}(V)$. //per-channel.
@@
     Load $\hat{Q}_{i}$ and $\delta_{Q}=\delta_{Q}[c_{w}*8i:c_{w}*8(i+1)]$ into an SM ;
@@
        Load $\hat{K}_{j}$, $\hat{V}_{j}$, and $\delta_{K}=\delta_{K}[4j:4(j+1)]$ into the SM ;
@@
        $S_{i}^{j}[st:st+c_{w}]=\psi^{-1}_{\delta_{Q}\delta_{K}}(\mathrm{Matmul}(\hat{Q}_{i}[st:st+c_{w}],\hat{K}_{j}^{\top}))+\Delta S_{j}$; // Paralleled by $c_{w}$ warps. The dequantization $\psi^{-1}_{\delta_{Q}\delta_{K}}$ is illustrated in Figure 6.
@@
        $O_{i}^{j}$(FP22) = $\mathrm{Matmul}((\widetilde{P}_{i}^{j}*448)\text{.to(FP8.e4m3)},V_{j})$ ;
        $O_{i}^{j}$(FP32) = $\mathrm{diag}(e^{m_{i}^{j-1}-m_{i}^{j}})^{-1}O_{i}^{j-1}(FP32)+O_{i}^{j}$(FP22) ;
@@
Table 4: Error of the FP8 Matmul instruction of mma(f8f8f32).
Precision of Accumulated Value E8M13 E8M23
3.4 FP32 MMA Buffer for FP22 Accumulator
To avoid the drawback of SageAttention where FP16 Matmul with an FP16 accumulator for $PV$ is only effective on GPUs such as the RTX4090, we adopt the approach of quantizing $P$ and $V$ to FP8 to leverage the universal acceleration of FP8 Tensor cores. However, we discover that the accumulator for the mma(f32f8f8f32) instruction on the Ada architecture is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. Specifically, for mma(f32f8f8f32) instruction $C=AB+D$, where $A,B$ are tensors in FP8 data type and $C,D$ are tensors with FP32 data type, we initialize the $A,B$ to zero and vary $D$ to test the data type of the accumulator. As shown in Table 4, when $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ matches the result of the mma(f16f16f32) instruction. However, when $D$ is initialized with more than 13 mantissa bits, the error of $C$ corresponds to the difference between the results of mma(f32f16f16f32) and mma(f32f8f8f32). Consequently, the matrix multiplication of $\widetilde{P}V$, quantized to FP8, incurs a certain degree of accuracy loss compared to using an FP32 accumulator.
To mitigate this accuracy loss as much as possible, we propose to use an FP32 buffer to accumulate the values of $\widetilde{P}_{i}^{j}V_{j}$ in FP22. Specifically, as shown in Algorithm 1, after each Matmul between the tiled $\widetilde{P}$ and $V$, we store the Matmul result in the FP32 buffer. This approach prevents the FP22 accumulator from accumulating errors across the entire sequence, confining errors to the tiled range instead.
@@ 3.4 FP32 MMA Buffer for FP22 Accumulator
Table 7: Two kernel implementations of SageAttention2.
      Kernel       $\psi_{Q}(Q)$, $\psi_{K}(K)$       $\psi_{P}(\widetilde{P})$, $\psi_{V}(V)$
      SageAttn2-4b       INT4, per-thread       FP8, per-block;  FP8, per-channel
      SageAttn2-8b       INT8, per-thread       FP8, per-block;  FP8, per-channel
3.5 Quantization for Q, K, P, V
Quantization for $Q,K$. We propose $\psi_{Q}(Q)$ and $\psi_{K}(K)$ in granularity of per-thread. This is first because per-channel quantization is not feasible since the scale factors of the inner axis of $QK^{\top}$ cannot used to do dequantization (Xiao et al., 2023a). Second, as shown in Table 3 and Table 3, we compare the average and worst-case accuracies of INT4 quantization at per-token, per-thread, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX. Results indicate that the accuracy of per-thread quantization is very close to per-token and outperforms a lot more than per-block and per-tensor. Furthermore, per-thread quantization incurs less dequantization overhead than per-token, as discussed in Section 3.3.
Quantization for $\widetilde{P},V$. We choose FP8, specifically the E4M3 data type, for $\psi_{P}(\widetilde{P})$ and $\psi_{V}(V)$ for two reasons. First, most GPUs have tensor cores that support FP8 Matmul operations, which are twice as fast as those using FP16. Second, Table 6 and Table 6 show the average and worst accuracy of different data types used for $\widetilde{P},V$ using real $Q,K,V$ across all layers of CogvideoX. We can see that the accuracy of using E4M3 is very close to that of using FP16 and superior to that of E5M2 and INT8. We propose to use $\psi_{P}(\widetilde{P})$ in per-block and $\psi_{V}(V)$ in per-channel for three reasons. First, per-channel quantization for $\widetilde{P}$ and per-token quantization for $V$ are not viable because dequantization requires scale factors of outer axis. Second, $\widetilde{P}=\mathrm{exp}(S_{i}-\mathrm{rowmax}(S_{i}))$, where $S_{i}$ is the Matmul result of a block of $Q$ and $K^{\top}$, the max value in each row of $\widetilde{P}$ is 1. Hence, we can assign a single static scale $s=\frac{1}{448}$ to a block $\widetilde{P}$. Third, per-channel quantization can address the channel-wised outlier of $V$.
@@ 3.5 Quantization for Q, K, P, V
3.6 An Optional Technique: Smooth V
We also figure out another way to mitigate the accuracy loss due to the FP22 accumulator when $V$ possesses channel-wise biases:
@@ 3.6 An Optional Technique: Smooth V
As shown in Figure 7, this strategy could enhance the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ consistently features channel-wise biases that are exclusively positive or negative, for instance, ranging between 8 and 9. Consequently, the values of $\widetilde{P}V$ could be quite large. However, the floating-point number representation range is not uniform—it is denser near zero. Therefore, by subtracting the mean along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in higher representational precision. We believe such an explanation and strategy could address lots of issues [^1] https://github.com/triton-lang/triton/issues/4476, https://github.com/triton-lang/triton/issues/5065 reported by the community. Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$. This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$. In other words, this method decomposes $V$ into two parts: $\overrightarrow{V_{m}}$ and $V$. For $V$, it centers the values of each column around zero, which leads to the dot product result between a row from the quantized $\widetilde{P}$ matrix and a column from the quantized $V$ matrix being closer to zero. This makes the representation of FP22 more accurate. Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.
Remark. This technique is optional and not employed in our main experiments. This is because its benefits are marginal, only significantly enhancing attention accuracy primarily when $V$ exhibits channel-wise outliers, which are absent in some models, such as $V$ in Llama3.1 (see Figure 4).
@@ 4.1 Setup
Models. We validate the effectiveness of SageAttention2 across a diverse set of representative models from language, image, and video generation. Specifically, we conduct experiments on ten models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B), CogvideoX (1.5-5B) (Yang et al., 2024), HunyuanVideo (Weijie Kong & Jie Jiang, 2024), and Mochi (Team, 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) and Stable-Diffusion3.5 (turbo) (Stability AI, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
Datasets. Text-to-text models are evaluated on four zero-shot tasks: WikiText (Merity et al., 2022) to assess the model’s prediction confidence, LAMBADA (Paperno et al., 2016) evaluate contextual understanding, MMLU (Hendrycks et al., 2020) for measuring knowledge across various subjects, and Longbench (Bai et al., 2023) for evaluating bilingual, multitask, and comprehensive assessment of long context understanding capabilities. Text-to-video models are evaluated using the open-sora (Zheng et al., 2024) prompt sets. Text-to-image models are assessed on MJHQ-30K (Li et al., 2024). TIMM is evaluated on on three image datasets: ImageNet (Deng et al., 2009), ImageNet-Sketch (Sketch) (Wang et al., 2019), and ImageNet-Rendition (ImageNet-r) (Hendrycks et al., 2021).
Metrics. For text-to-text models, we use perplexity (ppl.) (Jelinek et al., 1977) for WikiText, Accuracy (Acc.) for LAMBADA and MMLU, and Longbench score (Bai et al., 2023). For text-to-video models, following Zhao et al. (2024), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024) to measure the text-video alignment; (VQA-a) and (VQA-t) to assess the video aesthetic and technical quality, respectively; and Flow-score (FScore) for temporal consistency (Wu et al., 2023). For text-to-image models, generated images are compared with the images in MJHQ-30K dataset in three aspects: FID (Heusel et al., 2017) and sFID (Salimans et al., 2016) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2024) for human preference. For TIMM, we use classification accuracy.
@@
Baselines. (1) SmoothAttn. Following Qserve (Lin et al., 2024), we apply smooth quant for matrixes $Q,K$ with smoothing factor $\alpha=0.5$. (2) HadmdAttn. Following Quarot (Ashkboos et al., 2024), we apply random Hadamard transformation for matrixes $Q,K$ before INT4 quantization. (3) SageAttention (Zhang et al., 2024) uses smoothing $K$, INT8 per-block quantization for $Q,K$, and FP16 for $\widetilde{P},V$.
Figure 12: Comparison examples from HunyuanVideo, prompts are sampled from open-sora prompt sets.
Figure 13: Comparison examples from CogvideoX (2B), prompts are sampled from open-sora prompt sets.
Table 8: Average accuracy across all layers of CogvideoX using different smooth methods.
@@
Table 9: Worst accuracy across all layers of CogvideoX using different smooth methods.
@@
Table 10: End-to-end metrics across text and video generation models. - means the model does not support this dataset evaluation. ✗ indicates an inability to generate results for evaluation.
@@
HadmdAttn 6.771 0.860 0.360 -
SmoothAttn 6.717 0.867 0.392 -
SageAttn2-4b 5.912 0.881 0.428 -
SageAttn2-8b 5.828 0.886 0.438 -
SageAttention 5.824 0.887 0.439 -
@@
HadmdAttn 7.872 0.762 0.500 44.07
SmoothAttn 7.180 0.783 0.541 44.69
SageAttn2-4b 6.256 0.798 0.607 48.79
SageAttn2-8b 6.019 0.811 0.634 49.59
SageAttention 6.017 0.812 0.634 49.55
@@
HadmdAttn 7.989 0.435 0.669 45.97
SmoothAttn 8.943 0.449 0.592 42.20
SageAttn2-4b 7.352 0.433 0.725 49.23
SageAttn2-8b 7.242 0.432 0.745 49.60
SageAttention 7.243 0.433 0.744 49.79
@@
CogvideoX (2B) Full-Precision 0.1836 0.9975 77.605 75.360 3.006
HadmdAttn 0.1742 0.9877 29.780 23.985 0.499
SmoothAttn 0.1741 0.9870 41.703 47.043 0.624
SageAttn2-4b 0.1821 0.9973 77.368 74.906 2.603
SageAttn2-8b 0.1829 0.9977 76.532 74.281 2.941
SageAttention 0.1833 0.9976 76.997 71.360 2.988
CogvideoX (1.5-5B) Full-Precision 0.1778 0.9979 70.231 70.928 2.507
HadmdAttn 0.1576 0.9933 8.990 2.299
SmoothAttn 0.1559 0.9950 8.812 2.277
SageAttn2-4b 0.1721 0.9978 57.729 52.989 2.884
SageAttn2-8b 0.1775 0.9980 69.492 74.415 2.487
SageAttention
HunyuanVideo Full-Precision 0.1783 0.9995 82.516 75.934 0.604
HadmdAttn 0.1727 0.9989 7.514 0.762 0.175
SmoothAttn 0.1739 0.9988 6.987 0.609 0.148
SageAttn2-4b 0.1751 0.9995 81.478 65.371 0.610
SageAttn2-8b 0.1782 0.9996 81.786 75.354 0.586
SageAttention 0.1786 0.9995 82.496 79.843 0.597
Mochi Full-Precision 0.1798 0.9986 45.549 65.416 1.266
HadmdAttn 0.1733 0.9980 9.053 25.133 0.704
SmoothAttn 0.1687 0.9978 3.383 3.480 0.241
SageAttn2-4b 0.1783 0.9986 35.955 43.735 1.137
SageAttn2-8b 0.1797 0.9986 46.760 64.901 1.255
SageAttention 0.1800 0.9987 48.707 63.763 1.269
Table 11: Table 10 continued. End-to-end metrics across image generation and classification models.
@@
        Flux         Full-Precision        10.960        16.648        26.180        1.009
        HadmdAttn        11.353        18.495        26.123        0.965
        SmoothAttn        11.149        19.017        26.109        0.959
        SageAttn2-4b        10.577        17.497        26.141        0.998
        SageAttn2-8b        10.927        16.723        26.175        1.009
        SageAttention        10.944        16.641        26.171        1.008
               Stable-Dif        fusion3.5         Full-Precision        14.105        15.646        25.505        0.902
        HadmdAttn        14.259        15.909        25.513        0.886
        SmoothAttn        14.161        15.649        25.510        0.887
        SageAttn2-4b        14.097        15.397        25.487        0.895
        SageAttn2-8b        14.106        15.647        25.499        0.901
        SageAttention        14.140        15.678        25.503        0.902
@@
HadmdAttn 84.50% 44.89% 58.80%
SmoothAttn 84.40% 44.68% 58.73%
SageAttn2-4b 86.67% 45.24% 59.29%
SageAttn2-8b 84.79% 45.39% 59.57%
SageAttention 84.74% 45.38% 59.95%